# Quantum property testing in sparse directed graphs

### v1

@@
Simon Apers Research supported in part by the European QuantERA project QOPT (ERA-NET Cofund 2022-25), the French PEPR integrated projects EPiQ (ANR-22-PETQ-0007) and HQI (ANR-22-PNCQ-0002), and the French ANR project QUOPS (ANR-22-CE47-0003-01). Université Paris Cité, CNRS, IRIF, Paris, France Frédéric Magniez11footnotemark: 1 Université Paris Cité, CNRS, IRIF, Paris, France Sayantan Sen Research supported by the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme NRF2021-QEP2-02-P05. Centre for Quantum Technologies, National University of Singapore, Singapore Dániel Szabó11footnotemark: 1 Université Paris Cité, CNRS, IRIF, Paris, France
We initiate the study of quantum property testing in sparse directed graphs, and more particularly in the unidirectional model, where the algorithm is allowed to query only the outgoing edges of a vertex.
In the classical unidirectional model the problem of testing $k$-star-freeness, and more generally $k$-source-subgraph-freeness, is almost maximally hard for large $k$. We prove that this problem has almost quadratic advantage in the quantum setting. Moreover, we prove that this advantage is nearly tight, by showing a quantum lower bound using the method of dual polynomials on an intermediate problem for a new, property testing version of the $k$-collision problem that was not studied before.
@@ 1.1 Context
Graphs are of paramount importance for instance when it comes to understanding large datasets, since they provide a natural way to represent and analyze complex relationships inside datasets. Goldreich, Goldwasser, and Ron [goldreich1998property] were the first to consider graphs in the context of property testing. Formally, given some form of query access to an unknown graph $G$ on $N$ vertices, and a property $\mathcal{P}$ of interest, the goal is to distinguish with high probability if $G$ satisfies the property $\mathcal{P}$, or whether it is “far” from all graphs that satisfy $\mathcal{P}$, with a suitable notion of farness. In [goldreich1998property] the “dense” model was considered, where a graph is accessed through adjacency queries: for a pair of vertices $(u,v)$, the query reveals whether $(u,v)$ is an edge in the graph. In this model, a graph $G$ is $\varepsilon$-far from satisfying $\mathcal{P}$ if one needs to add or remove at least $\varepsilon N^{2}$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$.
In a later work, Goldreich and Ron [goldreich2002property] introduced the “bounded-degree” model for testing sparse graphs, focusing on the properties of bipartiteness and expansion. In this model, a $d$-bounded degree graph $G$ with $N$ vertices is accessed by performing neighbor queries: for a vertex $v$ and an integer $i\in[d]$, the query $(v,i)$ returns either the $i$-th neighbor of $v$, or some special symbol if $v$ has less than $i$ neighbors. The graph $G$ is said to be $\varepsilon$-far from some property $\mathcal{P}$, if one needs to add or delete at least $\varepsilon dN$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$. Over the last two decades, there has been a significant number of works in this model, and we refer the interested reader to the survey by Goldreich [goldreich2010introduction].
@@ 1.2 Property testing of directed bounded degree graphs
While all of the aforementioned works consider undirected graphs, many real-world instances (such as the world wide web) actually correspond to directed graphs. Consequently, Bender and Ron [bender2002testing] introduced a model of property testing for directed graphs, focusing on the properties of acyclicity and connectivity. Following that work, we open a new research line by studying quantum algorithms for testing directed graphs. As we will see, by doing so we address new fundamental questions in the field of quantum complexity. Answering them requires using recent techniques and partially answering some new or open questions.
@@
As described in [bender2002testing], for bounded-degree directed graphs there are two natural query models: (i) the unidirectional model, where the algorithm is allowed to query the outgoing edges of a vertex, but not the incoming edges, and (ii) the bidirectional model, where the algorithm can query both the incoming and outgoing edges of a vertex. Interestingly, [bender2002testing] showed that strong connectivity is testable in the bidirectional model (i.e., it can be tested with a number of queries that depends on $\varepsilon$ but not on $N$), but it requires $\Omega(\sqrt{N})$ queries in the unidirectional model. Later, the testability of other graph properties like Eulerianity, vertex and edge connectivity [orenstein2011testing, yoshida2010testing, forster2020computing, chen2019testability] was also shown in the bidirectional model. While there is a clear distinction between the two models, Czumaj, Peng and Sohler [czumaj2016relating] showed that if a property is testable in the bidirectional model, then it has a sublinear (i.e., $o(N)$) query complexity in the unidirectional model.
@@
In this work we consider a particularly important problem in the unidirectional model: the problem of testing subgraph-freeness. More precisely, we examine the problem of testing “$k$-source-subgraph-freeness”, where the goal is to test $H$-freeness for some constant-sized subgraph $H$ with $k$ “source components”, where a source component is a strongly connected subgraph that has no incoming edges. This problem was first studied by Hellweg and Sohler [hellweg2013property], and they provided a testing algorithm that performs $O(N^{1-1/k})$ queries. They also proved a tight lower bound of $\Omega(N^{2/3})$ for the $k=3$ case (see [hellweg2013property, Theorem 1 and Theorem 3]). Very recently, Peng and Wang [peng2023optimal] proved a matching lower bound for any constant $k$. In particular, they showed that $\Omega(N^{1-\frac{1}{k}})$ queries are necessary for testing $k$-star-freeness in the unidirectional model, for arbitrary $k$ (see [peng2023optimal, Theorem 1.2]). Notice that asymptotically the complexity of testing $k$-star-freeness becomes $\Omega(N)$. This also proves that the aforementioned reduction of [czumaj2016relating] cannot be made much stronger: for the property of $k$-star-freeness, the separation between the query complexities in the bi- and unidirectional models is maximal, because this property can be tested using constant number of queries in the bidirectional model.
1.3 Our results and techniques
@@
On the one hand, we prove that the problem of testing $k$-source-subgraph-freeness in the unidirectional model, which is almost maximally hard for large $k$ in the classical regime, has almost quadratic advantage in the quantum setting. Moreover, we prove that this quantum advantage is nearly tight, by showing a quantum lower bound using the method of dual polynomials.
On the other hand, we show that not all problems in graph property testing admit such a quantum speedup such as the problem of 3-colorability in the related undirected bounded-degree model, which requires $\Omega(N)$ queries in both classical and quantum models.
We start by describing the upper bound result for testing $k$-source-subgraph-freeness.
@@
In order to prove the above result, we connect it to the problem of finding $k$-collisions, which is one of the most-studied problems in quantum query complexity – see [zhandry_collision] and the references therein. In that work, they also prove a matching lower bound showing that their algorithm for finding $k$-collisions is optimal. However, the property testing variant of this problem – which is what we study – could be easier, and their lower bound technique does not apply to this case. This is why we use a different method to prove a lower bound, which is almost matching.
@@
We apply the well-studied dual polynomial method for proving this lower bound. We would like to point out that [ben2020symmetries] also used dual polynomial method for proving quantum lower bounds for other problems. However, their techniques do not directly apply for our purpose, and we make adaptations of them. Here, we consider the auxiliary problem of $k$-collision-freeness that we reduce to $k$-source-subgraph-freeness. In general, testing collision-freeness of functions is a difficult problem. In fact, both the classical works of [hellweg2013property] and [peng2023optimal] use a collision testing result from [raskhodnikova2009strong]. Moreover, the authors in [ambainis2016efficient] stated it as an open question if one could use a variant of the proportional moments technique of [raskhodnikova2009strong] for proving better quantum lower bounds. However, it is not clear if we can use the result of [raskhodnikova2009strong] directly for our purpose. Instead, we adapt the dual polynomial method to our context, which we believe will be of independent interest for improving the quantum lower bounds of other problems.
Doing so, we make the first step towards extending the techniques of [raskhodnikova2009strong]. At the heart of their lower bound is a construction of two positive integer random variables, $X_{1}$ and $X_{2}$, with different expectations but with the following conditions on the first $k-1$ moments: $\mathop{\mathbb{E}}[X_{1}]/\mathop{\mathbb{E}}[X_{2}]=\mathop{\mathbb{E}}[X^{2}_{1}]/\mathop{\mathbb{E}}[X^{2}_{2}]=\ldots=\mathop{\mathbb{E}}[X^{k-1}_{1}]/\mathop{\mathbb{E}}[X^{k-1}_{2}]$. Such a construction leads then to a randomized query complexity lower bound of $\Omega(N^{1-\frac{1}{k}})$ for various distinguishing problems such as $k$-collision-freeness [peng2023optimal]. We conjecture that a similar result holds in the quantum setting with a lower bound of $\Omega\Big{(}N^{\frac{1}{2}\left(1-\frac{1}{k}\right)}\Big{)}$. One can consider this work as a proof of this conjecture for the special case of $k$-collision-freeness, and we hope that it will serve as a step towards proving it in general.
Finally, even for the case of undirected graphs, we show that not all problems in graph property testing have a quantum advantage. For this we consider the property testing variant of the famous problem of $3$-colorability: namely, distinguishing whether the undirected graph $G$ can be properly colored with $3$ colors, or one needs to modify a large fraction of its edges to make it $3$-colorable. In the classical bounded degree setting, this problem has been studied by [bogdanov2002lower], who proved a lower bound of $\Omega(N)$ queries. In this work, we present a simple argument that proves that there exists no sublinear quantum tester either for this problem. Our result is stated as follows:

### v2

@@
Simon Apers Email: {apers, magniez, szabo}@irif.fr. Research supported in part by the ERC Advanced Grant PARQ (grant agreement No 885394), the European QuantERA project QOPT (ERA-NET Cofund 2022-25), the French PEPR integrated projects EPiQ (ANR-22-PETQ-0007) and HQI (ANR-22-PNCQ-0002), and the French ANR project QUOPS (ANR-22-CE47-0003-01). Université Paris Cité, CNRS, IRIF, Paris, France Frédéric Magniez11footnotemark: 1 Université Paris Cité, CNRS, IRIF, Paris, France Sayantan Sen Email: sayantan789@gmail.com. Research supported by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009). Centre for Quantum Technologies, National University of Singapore, Singapore Dániel Szabó11footnotemark: 1 Université Paris Cité, CNRS, IRIF, Paris, France
We initiate the study of quantum property testing in sparse directed graphs, and more particularly in the unidirectional model, where the algorithm is allowed to query only the outgoing edges of a vertex. In the classical unidirectional model, the problem of testing $k$-star-freeness, and more generally $k$-source-subgraph-freeness, is almost maximally hard for large $k$. We prove that this problem has almost quadratic advantage in the quantum setting. Moreover, we show that this advantage is nearly tight, by showing a quantum lower bound using the method of dual polynomials on an intermediate problem for a new, property testing version of the $k$-collision problem that was not studied before.
@@ 1.1 Context
Graphs are of paramount importance for instance when it comes to understanding large datasets, since they provide a natural way to represent and analyze complex relationships inside datasets. Goldreich, Goldwasser, and Ron [goldreich1998property] were the first to consider graphs in the context of property testing. Formally, given some form of query access to an unknown graph $G$ on $N$ vertices, and a property $\mathcal{P}$ of interest, the goal is to distinguish with high probability if $G$ satisfies the property $\mathcal{P}$, or whether it is “far” from all graphs that satisfy $\mathcal{P}$, with a suitable notion of farness. In [goldreich1998property], the “dense” model was considered, where a graph is accessed through adjacency queries: for a pair of vertices $(u,v)$, the query reveals whether $(u,v)$ is an edge in the graph. In this model, a graph $G$ is $\varepsilon$-far from satisfying $\mathcal{P}$ if one needs to add or remove at least $\varepsilon N^{2}$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$.
In a later work, Goldreich and Ron [goldreich2002property] introduced the “bounded-degree” model for testing sparse graphs, focusing on the properties of bipartiteness and expansion. In this model, a $d$-bounded degree graph $G$ with $N$ vertices is accessed by performing neighbor queries: for a vertex $v$ and an integer $i\in[d]$, the query $(v,i)$ returns either the $i$-th neighbor of $v$, or some special symbol if $v$ has less than $i$ neighbors. The graph $G$ is said to be $\varepsilon$-far from some property $\mathcal{P}$, if one needs to add or delete at least $\varepsilon dN$ edges of $G$ to obtain a graph that satisfies $\mathcal{P}$. Over the last two decades, there has been a significant number of works in this model, and we refer the interested reader to the books by Goldreich [goldreich2017introduction] and Bhattacharyya and Yoshida [bhattacharyya2022property] and several surveys [DBLP:journals/eatcs/Fischer01, DBLP:journals/fttcs/Ron09, Czumaj2010, DBLP:journals/siamdm/RubinfeldS11].
@@ 1.2 Property testing of directed bounded degree graphs
While all of the aforementioned works consider undirected graphs, many real-world instances (such as the world wide web) actually correspond to directed graphs. Consequently, Bender and Ron [bender2002testing] introduced a model of property testing for directed graphs, focusing on the properties of acyclicity and connectivity. Following that work, we open a new research line by studying quantum algorithms for testing directed graphs. As we will see, by doing so, we address new fundamental questions in the field of quantum complexity. Answering them requires using recent techniques and partially answering some new or open questions.
@@
As described in [bender2002testing], for bounded-degree directed graphs there are two natural query models: (i) the unidirectional model, where the algorithm is allowed to query the outgoing edges of a vertex, but not the incoming edges, and (ii) the bidirectional model, where the algorithm can query both the incoming and outgoing edges of a vertex. Interestingly, [bender2002testing] showed that strong connectivity is testable in the bidirectional model (i.e., it can be tested with a number of queries that depends on $\varepsilon$ but not on $N$), but it requires $\Omega(\sqrt{N})$ queries in the unidirectional model. Later, the testability of other graph properties like Eulerianity, vertex and edge connectivity [orenstein2011testing, yoshida2010testing, forster2020computing, chen2019testability] was also shown in the bidirectional model. While there is a clear distinction between the two models, Czumaj, Peng and Sohler [czumaj2016relating] showed that if a property is testable in the bidirectional model, then it has a tester with sublinear (i.e., $o(N)$) query complexity in the unidirectional model.
@@
In this work, we consider a particularly important problem in the unidirectional model: the problem of testing subgraph-freeness. More precisely, we examine the problem of testing “$k$-source-subgraph-freeness”, where the goal is to test $H$-freeness for some constant-sized subgraph $H$ with $k$ “source components”, where a source component is a strongly connected subgraph that has no incoming edges. This problem was first studied by Hellweg and Sohler [hellweg2013property], and they provided a testing algorithm that performs $O(N^{1-1/k})$ queries. They also proved a tight lower bound of $\Omega(N^{2/3})$ for the $k=3$ case (see [hellweg2013property, Theorem 1 and Theorem 3]). Very recently, Peng and Wang [peng2023optimal] proved a matching lower bound for any constant $k$. In particular, they showed that $\Omega(N^{1-\frac{1}{k}})$ queries are necessary for testing $k$-star-freeness (which is a special case of testing $k$-source-subgraph-freeness) in the unidirectional model, for arbitrary $k$ (see [peng2023optimal, Theorem 1.2]). Notice that asymptotically the complexity of testing $k$-star-freeness becomes $\Omega(N)$. This also proves that the aforementioned reduction of [czumaj2016relating] can not be made much stronger: for the property of $k$-star-freeness, the separation between the query complexities in the bi- and unidirectional models is maximal, as this property can be tested using constant many queries in the bidirectional model.
1.3 Related works on collision finding
A closely related problem to finding $k$-stars in graphs is finding $k$-collisions in integer sequences. The two mentioned classical papers on subgraph-freeness testing [hellweg2013property, peng2023optimal] actually consider a collision-type intermediate problem for proving their lower bounds. As we are also going to do so, let us look at some related, known results.
The problem of collision finding is a ubiquitous problem in the field of algorithm theory with wide applications in cryptography. Here, given a sequence $s$ of $N$ integers, the goal is to find a duplicate in $s$. If one has the guarantee that $\Theta(N)$ elements of the sequence are duplicated, which is the case for instance when the sequence consists of uniformly random integers from $[N]$, it is well-known that classically $\Theta(\sqrt{N})$ queries are necessary and sufficient due to the birthday paradox. In the quantum model, this can be solved with query complexity $\Theta(N^{1/3})$ by the algorithm of Brassard, Høyer and Tapp [BHT]. The matching lower bound was first stated for a specific set of hard instances known as $2$-to-$1$ (i.e., each integer appears exactly twice or not at all) by Aaronson and Shi [AaronsonShi]. For some constant integer $k\geq 3$, those results can be further extended to finding $k$-collisions in a random input with suitable alphabet size, so that it contains $\Theta(N)$ $k$-duplicates with high probability. The classical query complexity for this problem is $\Theta(N^{1-1/k})$ [hellweg2013property, peng2023optimal], and quantumly it is $\Theta\left(N^{\frac{1}{2}\left(1-\frac{1}{2^{k}-1}\right)}\right)$ [zhandry_collision]. The situation is more complex for non-random inputs.
Remarkably, the complexity of testing $k$-collision-freeness (i.e., the absence of $k$-collisions) is harder to settle on the lower bound side than the finding version. In this work, we are going to focus on the hardness of distinguishing inputs that have linearly many collisions from those that do not have any. For $k=2$, the two problems have the same complexity, since intuitively the only way to distinguish is to find a collision. This can be formalized easily in the classical case. Quantumly, this is more challenging, but the lower bound in [AaronsonShi] proved the hardness of distinguishing between $2$-to-$1$ instances and ones with no duplicate.
However, for larger $k$, distinguishing such inputs might be easier than finding a collision. The classical upper bound of $O(N^{1-1/k})$ queries is straightforward for the finding variant. In the lower bounds of [hellweg2013property, peng2023optimal], the authors consider the distinguishing version, so classically the question is settled. But in the quantum setting, the upper and lower bounds of [zhandry_collision] are tight only for finding $k$-collisions in random inputs, and for the distinguishing variant, we are not currently aware of anything better than the $\Omega(N^{1/3})$ lower bound corresponding to the $k=2$ case. To our knowledge, this problem has not yet been studied in the quantum setting.
1.4 Our results
@@
In the first line, we consider the problem of testing $k$-source-subgraph-freeness in the unidirectional model. This problem is almost maximally hard for large $k$ in the classical regime, and we show that it admits an almost quadratic advantage in the quantum setting.
@@
In order to prove the above result, we connect it to the problem of finding $k$-collisions. In [zhandry_collision], an algorithm is given for finding $k$-collisions in sequences of random integers. We generalize this to the context of graph property testing in two ways: first, finding a subgraph (instead of a collision); and second, considering graphs that are far from being $H$-free (instead of random).
Moreover, we prove that this quantum advantage is nearly tight, by showing a quantum lower bound using the method of dual polynomials.
@@
For proving graph property testing lower bounds, both the classical works of [hellweg2013property] and [peng2023optimal] prove collision testing lower bounds using the proportional moments technique of [raskhodnikova2009strong]. At the heart of this technique is a construction of two positive integer random variables, $X_{1}$ and $X_{2}$, with different expectations but with the following conditions on the first $k-1$ moments: $\mathop{\mathbb{E}}[X_{1}]/\mathop{\mathbb{E}}[X_{2}]=\mathop{\mathbb{E}}[X^{2}_{1}]/\mathop{\mathbb{E}}[X^{2}_{2}]=\ldots=\mathop{\mathbb{E}}[X^{k-1}_{1}]/\mathop{\mathbb{E}}[X^{k-1}_{2}]$. Such a construction leads to a randomized query complexity lower bound of $\Omega(N^{1-\frac{1}{k}})$ for various property testing problems such as $k$-collision-freeness [peng2023optimal]. Having a quantum version of this technique has been identified as an important open problem [ambainis2016efficient], since this could be used to pave the way to stronger quantum lower bounds in related settings. We modestly made progress to this quest for the special case of testing $k$-collision-freeness.
In [zhandry_collision], in addition to the algorithm we mentioned, they also prove a matching lower bound showing that their algorithm for finding $k$-collisions in random inputs is optimal. However, this time we cannot reuse those techniques for our purpose for two main reasons. First, the property testing variant of this problem could be easier. Moreover, their lower bound technique requires random inputs and hence it does not apply to our case. This is why we use yet another method, that of dual polynomials, to prove our lower bound.
Theorem 1.3 (Restated in Theorem 4.1).
In the second line of results, we show that not all problems in graph property testing admit such a quantum speedup. This fact even remains valid for the case of undirected graphs. For this, we consider the property testing variant of the famous problem of $3$-colorability: namely, distinguishing whether an unknown undirected graph $G$ can be properly colored with $3$ colors, or one needs to modify a large fraction of its edges to make it $3$-colorable. In the classical bounded degree setting, this problem has been studied by [bogdanov2002lower], who proved a lower bound of $\Omega(N)$ queries. In this work, we present a simple argument that proves that there exists no sublinear quantum tester either for this problem. Our result is stated as follows: